OpenStudy (anonymous):

Find the linearization L(x) of y=e^(3x)ln(x) at a=1

OpenStudy (anonymous):

derivative of the function: e^(3x)(3xlog(x)+1)/x

OpenStudy (helder_edwin):

do u mean \[ y-y(1)=y'(1)(x-1) \] ?

OpenStudy (anonymous):

Yea I think thats what it mean, I think linerization is just another word for finding equation of the tangent line but I am not 100% sure

OpenStudy (helder_edwin):

u can aproximate any curve with its derivatives. the easiest would be the tangent line, as you said

OpenStudy (anonymous):

Yea what I did so far was use the equation you used, however when I typed in numbers I got them wrong so maybe I am solving it wrong?

OpenStudy (helder_edwin):

since \[ y=e^{3x}\ln x \] then \[ y(1)=e^3\ln 1=0 \] right?

OpenStudy (helder_edwin):

then \[ y'=3e^{3x}\ln x+e^{3x}\frac{1}{x}=e^{3x}\left(\ln x+\frac{1}{x}\right) \] then y'(1)=e^3(\ln 1+1)=e^3

OpenStudy (anonymous):

Yea, and then I got 50 for the answer of the derivative form one.

OpenStudy (helder_edwin):

\[ y'(1)=e^3(\ln 1+1)=e^3 \]

OpenStudy (helder_edwin):

so the tangent line would be \[\large y=e^3(x-1) \]

OpenStudy (anonymous):

ah I got it, for some reason I kept getting 50 for the one answer l(x)=20.08553692x-20.08553692 Thanks you!

OpenStudy (helder_edwin):

u r welcome